# The Fellowship of the LLMs: Multi-Agent Workflows for Synthetic Preference Optimization Dataset Generation

### v3

@@ 4.1 LLM-as-Evaluator
Table 2 shows the results of LLM-as-a-Judge approach on the three prompting strategies. The Independent Scoring prompt strategy consistently under-performs compared to the Direct Comparison and Combined Scoring approaches across all evaluated LLMs. This result is reflected in lower Cohen’s Kappa values in Table 2. In evaluating responses in isolation the LLM has to re-calibrate its scoring mechanism for every new response. This can lead to inconsistencies, especially when multiple responses are closely matched in quality. Due to the low Kappa values observed, we opted not to conduct experiments with the scoring-out-of-5 and 100 scales for Independent Scoring.
@@
The LLM-as-Judge evaluations, as shown in Table 2, indicate that GPT-4o outperforms all the models on PandaLM-Eval and MT-Bench achieving a Cohen’s Kappa score of 0.688 and 0.410 respectively. Additionally, GPT-4o consistently ranks in second position across all three datasets. This consistent top-tier performance underscores GPT’s effectiveness as a reliable judge in evaluating LLM responses. Gemma-2-27b outperforms all other models on the Fair-Eval dataset, achieving the highest score in this particular evaluation. However, it’s important to note that the Fair-Eval dataset is relatively small, consisting of only 80 samples. Furthermore, the Fair-Eval dataset primarily compares GPT-3.5-Turbo with Vicuna-13b, which might introduce a bias in favor of GPT models when GPT is the evaluator. Figure 4 shows that GPT-4o selects GPT-3.5-Turbo as the better agent 50 times and Vicuna-13b 30 times. This indicates a potential bias in favor of GPT responses when GPT-4o is the evaluator. Additionally, we can observe in the figure that Llama models also display a similar bias towards GPT responses, whereas Gemma models do not exhibit this bias, suggesting that Gemma is more impartial in its evaluations.
Fair-Eval PandaLM-Eval MT-Bench
@@
The LLM Debate approach, as summarized in Table 5, showcases varying degrees of effectiveness across three different datasets: Fair-Eval, PandaLM-Eval, and MT-Bench. GPT-4o performs the best across all datasets, with Cohen’s Kappa scores of 0.404, 0.654, and 0.402 respectively. LLM Debate outperforms LLM-as-a-Judge on Fair-Eval only and does not surpass the LLMs-as-a-Jury approach on any dataset. On Fair-Eval using the Debate framework increases the Kappa score of GPT-4o from 0.327 to 0.404 and of GPT-4o-mini from 0.333 to 0.360. It shows that the debate approach decreases the bias of GPT-4o and GPT-4o-mini towards the responses of it’s family.
Fair-Eval PandaLM-Eval MT-Bench
@@ 4.2 LLM-as-Generator
We compare the performance of Multi-Agent Feedback Loop with the baseline Single-Agents (GPT-4o, Llama-3.1-8b, Gemma-2-9b) using win rate as shown in Table 6. We utilize GPT-4o-as-a-judge in this evaluation process. For the baseline we find the win rate of Gemma and Llama against GPT-4o and each other. Both smaller models have similar win rate of 38.6% and 39.2% against GPT, while Gemma has a win rate of 66.6% against Llama.
@@ 4.3 Preference Optimization Dataset
Single-Agent Improvement.
For the Single-Agent improvement dataset generation we use GPT-4o-as-a-Judge in the evaluation module. In the generation module, we use LLM Feedback Loop with Llama-3.1-8b as the generator and Gemma-2-9b as the reviewer. The framework is shown in Figure 6. For the dataset generation, we use $N=3$ iterations. We pick the best response (judged by GPT-4o) as accepted and the other 2 responses as rejected and generate two datasets, one for DPO and one for KTO.
@@
Multi-Agent Improvement.
For the Multi-Agent improvement dataset generation we use Gemma-2-9b because the win rate for Gemma against Llama-3.1-8b is 66.6%. We give Gemma user prompt, human response and it’s own response and ask it to generate feedback based on the human reference. For the rejected responses we use the feedback without human reference. We generate two datasets, one for DPO and one for KTO.
Finally, we combine the two categories of dataset to make our final PO dataset to improve LLM for single-agent setting and LLM for multi-agent setting.
@@ 6 Future Work
7 Acknowledgement
@@ Appendix A System Prompts
Table 7 contains the three categories of system prompts tested for LLM-as-a-Judge approach. The winning prompt with Combined Scoring was used for LLMs-as-a-Jury. These prompts are modified versions of those used by (Zheng et al. 2023a). Table 8 present the system prompt and user message structure for LLM Debate and 9 shows the prompt for each role in the debate. This is based on the system prompt and the input structure used by (Chan et al. 2023). Table 10 shows the user message structure for the generator LLM and Table 11 shows the system prompt and user message for reviewer LLM in LLM Feedback Loop. Table 12 shows the system prompt and user message structure used for Multi-Agent improvement dataset generation.
@@ Appendix B Code and Datasets
The evaluation code, all the evaluation outputs and the generated datasets are publicly available on GitHub [^4] https://github.com/ulrs0/MA-PO. For evaluation of LLMs-as-Evaluators we used Alpaca-Eval [^5] https://huggingface.co/datasets/tatsu-lab/alpaca˙eval, Fair-Eval [^6] https://github.com/i-Eval/FairEval, PandaLM-Eval [^7] https://github.com/WeOpenML/PandaLM and MT-Bench [^8] https://huggingface.co/datasets/lmsys/mt˙bench˙human˙judgments. For evaluation of LLMs-as-Generators and single-agent improvement dataset generation we use the prompts from Argilla Capybara DPO dataset [^9] https://huggingface.co/datasets/argilla/distilabel-capybara-dpo-7k-binarized. For multi-agent improvement dataset generation we use prompts from No-Robots [^10] https://huggingface.co/datasets/HuggingFaceH4/no˙robots dataset.
@@ Appendix C Computing Infrastructure
We use the API for GPT-4o and GPT-4o-mini from OpenAI [^11] https://platform.openai.com/docs/overview. For Gemma and Llama models API from TogetherAI [^12] https://docs.together.ai/docs/introduction was used. We use Python3 libraries for both APIs and the temperature for the models was set to 0 for reproduciblity. For each evaluation, one run of the code was done.
Appendix D Reproducibility Checklist
  1. 1.
    Includes a conceptual outline and/or pseudocode description of AI methods introduced: Yes
  2. 2.
    Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results: Yes
  3. 3.
    Provides well marked pedagogical references for less-familiare readers to gain background necessary to replicate the paper: Yes
  4. 4.
    Does this paper make theoretical contributions: No
  5. 5.
    Does this paper rely on one or more datasets: Yes
    1. (a)
      A motivation is given for why the experiments are conducted on the selected datasets: Yes
    2. (b)
      All novel datasets introduced in this paper are included in a data appendix: Yes
    3. (c)
      All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes: Yes
    4. (d)
      All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations: Yes
    5. (e)
      All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available: Yes
    6. (f)
      All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisficing: NA
  6. 6.
    Does this paper include computational experiments: Yes
    1. (a)
      Any code required for pre-processing data is included in the appendix: Yes
    2. (b)
      All source code required for conducting and analyzing the experiments is included in a code appendix: Yes
    3. (c)
      All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes: Yes
    4. (d)
      All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from: Yes
    5. (e)
      If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results: Yes
    6. (f)
      This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks: Yes
    7. (g)
      This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics: Yes
    8. (h)
      This paper states the number of algorithm runs used to compute each reported result: Yes
    9. (i)
      Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information: No
    10. (j)
      The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank): Yes
    11. (k)
      This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments: Yes
    12. (l)
      This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting: NA
@@ Appendix D Reproducibility Checklist
Independent Scoring Please act as an impartial judge and evaluate the quality of the response provided by an AI assistants to the user question displayed below. Assign an overall score out of 10, where a higher score indicates better overall performance. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of their response. Begin your evaluation by comparing the two responses and provide a short explanation. Do not allow the length of the response to influence your evaluation.
@@
System Prompt We would like to request your feedback on the performance of two AI assistants in response to the user question. There are a few other referee assigned the same task, it’s your responsibility to discuss with them and think critically before you make your final judgement.
@@
General Public You are now General Public, one of the referees in this task. You are interested in the story and looking for updates on the investigation. Please think critically by yourself and note that it’s your responsibility to choose one of which is the better first.
Now it’s your turn to speak General Public, please make your talk short and clear.
@@
Psychologist You are now Psychologist, one of the referees in this task. You will study human behavior and mental processes in order to understand and explain human behavior. Please help other people to determine which response is the better one.
Now it’s your turn to speak Psychologist, please make your talk short and clear.
@@
Critic You are now Critic, one of the referees in this task. You will check fluent writing, clear sentences, and good wording in summary writing. Your job is to question others judgement to make sure their judgement is well-considered and offer an alternative solution if two responses are at the same level.
Now it’s your turn to speak Critic, please make your talk short and clear.
**Critic**:”
@@
Do not engage in formalities such as ’Thank you for your feedback’ or ’Here is an updated version…’ etc, just update the response.
User Message (Double Feedback) Update your response based on the feedback by the two assistant:
@@
Do not engage in formalities such as ’Thank you for your feedback’ or ’Here is an updated version…’ etc, just update the response.
@@
Table 12: The prompt and user message structure for the feedback generation for Multi-Agent improvement dataset.
Message Type Prompt
Your evaluation should consider factors such as the instruction following (the response should align with the user instructions), helpfulness, relevance, accuracy, and creativity of the response. Use the human answer as the reference answer, so that the evaluation and feedback is based on how to get closer to human answer.
Use the following format to respond:
### Evaluation:
[Add your evaluation here]
### Overall Score:
X/10
### Feedback:
[Add your feedback here]
User Message [Start of User Question]
{User Question}
[End of User Question]
[Start of Human’s Response]
{Human’s Response}
[End of Human’s Response]
[Start of Assistant’s Response]
{Assistant’s Response}
[End of Assistant’s Response]
Base your evaluation and feedback on human response. We want the LLM to mimic the human response. Do not mention the word ’human response’ in your evaluation and feedback. The user should not know that you have a reference answer.

### v4

@@ 4.1 LLM-as-Evaluator
Table 2 shows the results of LLM-as-a-Judge approach on the three prompting strategies.
@@
The Independent Scoring prompt strategy consistently under-performs compared to the Direct Comparison and Combined Scoring approaches across all evaluated LLMs. This result is reflected in lower Cohen’s Kappa values ranging from only 0.170 to 0.254 in Table 2. In evaluating responses in isolation the LLM has to re-calibrate its scoring mechanism for every new response. This can lead to inconsistencies, especially when multiple responses are closely matched in quality. Due to the low Kappa values observed, we opted not to conduct experiments with the scoring-out-of-5 and 100 scales for Independent Scoring.
@@
The LLM-as-Judge evaluations, as shown in Table 2, indicate that GPT-4o outperforms all the models on PandaLM-Eval and MT-Bench achieving a Cohen’s Kappa score of 0.688 and 0.410 respectively. Additionally, GPT-4o consistently ranks in second position across all three datasets. This consistent top-tier performance underscores GPT’s effectiveness as a reliable judge in evaluating LLM responses. Gemma-2-27b outperforms all other models on the Fair-Eval dataset, achieving the highest score in this particular evaluation. However, it’s important to note that the Fair-Eval dataset is relatively small, consisting of only 80 samples. Furthermore, the Fair-Eval dataset primarily compares GPT-3.5-Turbo with Vicuna-13b, which might introduce a bias in favor of GPT models when GPT is the evaluator.
Fair-Eval PandaLM MT-Bench
@@
Figure 3 shows that GPT-4o selects GPT-3.5-Turbo as the better agent 50 times and Vicuna-13b 30 times. This indicates a potential bias in favor of GPT responses when GPT-4o is the evaluator. Additionally, we can observe in the figure that Llama models also display a similar bias towards GPT responses, whereas Gemma models do not exhibit this bias, suggesting that Gemma is more impartial in its evaluations.
@@
The LLM Debate approach, as summarized in Table 5, showcases varying degrees of effectiveness across three different datasets: Fair-Eval, PandaLM-Eval, and MT-Bench.
Fair-Eval PandaLM MT-Bench
@@
GPT-4o performs the best across all datasets, with Cohen’s Kappa scores of 0.404, 0.654, and 0.402 respectively. LLM Debate outperforms LLM-as-a-Judge on Fair-Eval only and does not surpass the LLMs-as-a-Jury approach on any dataset. On Fair-Eval using the Debate framework increases the Kappa score of GPT-4o from 0.327 to 0.404 and of GPT-4o-mini from 0.333 to 0.360. It shows that the debate approach decreases the bias of GPT-4o and GPT-4o-mini towards the responses of it’s family.
@@ 4.2 LLM-as-Generator
We compare the performance of Multi-Agent Feedback Loop with the baseline Single-Agents (GPT-4o, Llama-3.1-8b, Gemma-2-9b) using win rate as shown in Table 6.
@@
We utilize GPT-4o-as-a-judge in this evaluation process. For the baseline we find the win rate of Gemma and Llama against GPT-4o and each other. Both smaller models have similar win rate of 38.6% and 39.2% against GPT, while Gemma has a win rate of 66.6% against Llama.
@@ 4.3 Preference Optimization Dataset
We use GPT-4o-as-a-Judge in the evaluation module because of its consistency and reliability as a judge across multiple datasets. In the generation module, we use LLM Feedback Loop with Llama-3.1-8b as the generator and Gemma-2-9b as the reviewer because of it’s highest win-rate against other configurations. The framework is shown in Figure 5. For the dataset generation, we use $N=3$ feedback iterations. For each prompt, we generate three responses using the generation module. These responses are then evaluated by GPT-4o in the evaluation module. The response judged to be the best by GPT-4o is labeled as accepted, while the other two responses are labeled as rejected to form the DPO and KTO datasets.
@@ 6 Future Work
7 Limitations
While our study demonstrates the potential of multi-agent workflows in automating the generation of PO datasets, several limitations should be acknowledged. Firstly, the use of multi-agent frameworks significantly increases computational complexity and resource consumption compared to single-agent models. The iterative processes in both the response generation and evaluation modules require more computational power and time, which may not be feasible for practitioners with limited resources. Additionally, GPT-4o is a proprietary model, which may not be accessible to all researchers, potentially hindering reproducibility and wider adoption of our methods.
8 Ethical Considerations
The automation of response evaluation and generation in PO dataset creation raises several ethical considerations that warrant careful attention. Relying on LLMs to simulate human judgments may perpetuate existing biases present in the training data of these models. If not properly addressed, this could result in PO datasets that reinforce stereotypes or unfairly represent certain groups, leading to biased behaviors in models fine-tuned on these datasets. The potential displacement of human annotators poses an ethical dilemma. While automation can increase efficiency and scalability, it may reduce opportunities for human involvement in the annotation process, affecting those who rely on such tasks for employment. Balancing automation with human oversight is essential to maintain ethical standards and ensure diverse perspectives are included.
In conclusion, while our approach offers advancements in automating PO dataset generation, it is imperative to remain vigilant about these ethical concerns. Implementing strategies to mitigate biases, maintaining transparency, involving human oversight, and adhering to ethical guidelines are essential steps in responsible AI development.
Acknowledgments
@@ Appendix A System Prompts
Table LABEL:tab:prompt-llm-judge contains the three categories of system prompts tested for LLM-as-a-Judge approach. The winning prompt with Combined Scoring was used for LLMs-as-a-Jury. These prompts are modified versions of those used by (Zheng et al., 2023a). Table LABEL:tab:prompt-llm-debate present the system prompt and user message structure for LLM Debate and LABEL:tab:prompt-roles shows the prompt for each role in the debate. This is based on the system prompt and the input structure used by (Chan et al., 2023). Table LABEL:tab:generator_prompt shows the user message structure for the generator LLM and Table LABEL:tab:reviewer_prompt shows the system prompt and user message for reviewer LLM in LLM Feedback Loop.
@@ Appendix B Code and Datasets
The evaluation code, all the evaluation outputs and the generated datasets are publicly available on GitHub [^4] https://github.com/ulrs0/MA-PO. For evaluation of LLMs-as-Evaluators we used Alpaca-Eval [^5] https://huggingface.co/datasets/tatsu-lab/alpaca_eval, Fair-Eval [^6] https://github.com/i-Eval/FairEval, PandaLM-Eval [^7] https://github.com/WeOpenML/PandaLM and MT-Bench [^8] https://huggingface.co/datasets/lmsys/mt_bench_human_judgments. For evaluation of LLMs-as-Generators and single-agent improvement dataset generation we use the prompts from Argilla Capybara DPO dataset [^9] https://huggingface.co/datasets/argilla/distilabel-capybara-dpo-7k-binarized. For multi-agent improvement dataset generation we use prompts from No-Robots [^10] https://huggingface.co/datasets/HuggingFaceH4/no_robots dataset. Alpaca-Eval and PandaLM-Eval are under Apache 2.0 license, Fair-Eval dataset is under CC BY 4.0 license, Argilla Capybara DPO is also under Apache 2.0 license. All datasets used in this paper comply with their respective license.
@@ Appendix C Computing Infrastructure
We use the API for GPT-4o and GPT-4o-mini from OpenAI [^11] https://platform.openai.com/docs/overview. For Gemma and Llama models API from TogetherAI [^12] https://docs.together.ai/docs/introduction was used. We use Python3 libraries for both APIs and the temperature for the models was set to 0 for reproduciblity. For each evaluation, one run of the code was done. OpenAI GPT-4o has a proprietary license. Llama-3.1 is under Llama-3.1 license and Gemma-2 is under Gemma license. All models used in this paper comply with their respective license.
@@
Independent Scoring Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Assign an overall score out of 10, where a higher score indicates better overall performance. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of their response. Begin your evaluation by comparing the two responses and provide a short explanation. Do not allow the length of the response to influence your evaluation.
@@
Table 7: The three types of system prompts for LLM-as-a-Judge and LLMs-as-a-Jury (continued).
@@
System Prompt We would like to request your feedback on the performance of two AI assistants in response to the user question. There are a few other referees assigned the same task; it’s your responsibility to discuss with them and think critically before you make your final judgement.
@@
Table 8: The system prompt and the user message structure for LLM Debate (continued).
@@
General Public You are now General Public, one of the referees in this task. You are interested in the story and looking for updates on the investigation. Please think critically by yourself and note that it’s your responsibility to choose which of the responses is better first.
Now it’s your turn to speak, General Public. Please make your talk short and clear.
@@
Psychologist You are now Psychologist, one of the referees in this task. You will study human behavior and mental processes in order to understand and explain human behavior. Please help others determine which response is the better one.
Now it’s your turn to speak, Psychologist. Please make your talk short and clear.
@@
Critic You are now Critic, one of the referees in this task. You will check for fluent writing, clear sentences, and good wording in summary writing. Your job is to question others’ judgment to make sure their judgment is well-considered and offer an alternative solution if two responses are at the same level.
Now it’s your turn to speak, Critic. Please make your talk short and clear.
**Critic**:
Table 9: The prompt for each role used in LLM Debate (continued).
@@
Do not engage in formalities such as ’Thank you for your feedback’ or ’Here is an updated version…’ etc., just update the response.
User Message (Double Feedback) Update your response based on the feedback by the two assistants:
@@
Do not engage in formalities such as ’Thank you for your feedback’ or ’Here is an updated version…’ etc., just update the response.
Table 10: The user message structure for the generator in LLM Feedback (continued).
@@
Table 11: The prompt and user message structure for the reviewer in LLM Feedback (continued).